I do apolgize. I just submitted a problem but I guess I didn't write it correctly. I'm still very new at this.
the question is
4^(x+1)=64 the x+1 are in exponential form. Where appropriate, include approximations to the nearest thousandth.
Again thank you so much.

Found 2 solutions by stanbon, khwang:
Answer by stanbon(57307)

I'm going to assume you do not want to solve this problem
using logarithms. If you do the procedure is quite different.
Without using logs you proceed as follows.
Rewrite the equation as follows:
4^(x+1) = 4^3
Then x+1 = 3
x = 4
Hope this helps.
Cheers,
Stan H.

Answer by khwang(438)

x=2 instead
and x^-y^2 = 53 is a hyperboa.
Kenny